# A theoretical analysis of mass scaling techniques

### v1

@@ 1 Introduction and background
Finite element analysis (FEA) has long established itself as an indispensable tool in many engineering disciples. Particularly in structural dynamics, one of its founding applications, where the method is extensively used for simulating the deformation and vibration of various structures and solids, including plates, shells and beams. In this context, explicit time integration is favored for various reasons. Firstly, the dynamics of the problem impose a natural restriction on the step size [1]. Enlarging it may lead to convergence issues within implicit unconditionally stable methods, specifically for nonlinear problems requiring the solution of a nonlinear system of equations at each iteration. On the contrary, the smaller step size used within explicit methods ensures greater robustness. Secondly, explicit methods offer the possibility of “lumping” the mass matrix enabling memory-efficient matrix-free implementations. Mass lumping consists in directly substituting the mass matrix with an ad hoc diagonal approximation [2, 3, 4, 5]. Thus, explicit methods may avoid solving costly linear or nonlinear systems of equations, left alone the burden of linearizing the stiffness.
@@
$\Delta t_{c}=\frac{2}{\sqrt{\lambda_{n}}}$ (1.1)
@@
where $\sqrt{\lambda_{n}}$ is the largest frequency of the discrete system [6, 7]. The largest discrete frequencies of classical $C^{0}$ finite element methods are wholly inaccurate and fortunately rarely contribute significantly to the solution. They form the so-called “optical branches” of the spectrum and diverge with the mesh size and polynomial degree [8, 9]. Smooth isogeometric analysis (IGA) [10] features far fewer inaccurate frequencies [11, 12, 13, 14]. They usually form noticeable spikes in the upper part of the spectrum and were coined outliers for this very reason [11]. Regardless of the discretization technique and smoothness, large inaccurate eigenvalues severely restrict the critical time step and removing or dampening them is paramount in explicit dynamics. The method for removing them, however, usually depends on the discretization technique. In isogeometric analysis, so-called outlier removal techniques commonly exploit the smoothness and tensor product nature of spline spaces (see e.g. [11, 14, 15, 16, 17, 18, 19]). These methods may also preserve the consistency of the discrete formulation and therefore higher rates of convergence. In more general cases with significantly less structure, one may resort to mass scaling. Broadly speaking, mass scaling consists in dampening higher frequencies by altering the discrete formulation and, more specifically, the mass matrix. Conventional mass scaling (CMS), which scales down all frequencies uniformly, is the most straightforward way of achieving this goal. Unfortunately, it also heavily deteriorates the accuracy of the lower frequencies and mode shapes, and, subsequently, the numerical solution. Selective mass scaling (SMS) is specifically designed for scaling down the highest frequencies while somewhat preserving the lowest ones. Thus, it still affects all eigenvalues, but not uniformly. Mathematically speaking, the idea is reminiscent of inexact deflation techniques, which consist in removing unwanted eigenvalues from the spectrum of a matrix. Exact deflation techniques have been known for decades and originated from the early work of Hotelling [20] for standard eigenproblems but are mostly impractical. Practical strategies, either global or local, are instead essentially heuristic and oftentimes little is known of their theoretical properties. Global strategies directly modify the assembled system matrices while local ones operate on the element matrices prior to assembly. The former are generally easier to analyze than the latter and are among the oldest methods. In the late 1990s, Macek and Aubert [21] and later Olovsson et al. [1] independently proposed the same global mass scaling procedure, based on linear fractional transformations of matrix pairs. The method was later generalized in [22] to rational polynomials, where the authors also stated optimality requirements for SMS. These methods are fully understood and transform (nonlinearly) the eigenvalues of a matrix pair while preserving its eigenvectors. Although the methods affect all eigenvalues, they may heavily damp the largest ones, while nearly preserving the smallest ones. In the same article, the authors also proposed an exact deflation method only scaling down the largest eigenvalues of the system. The idea was later revived and thoroughly improved in [23] but remains wholly impractical, unless only a few well-separated eigenvalues must be deflated [19]. As we will see, all of these methods are merely adaptions or reformulations of well-known linear algebra facts, mostly known since decades.
@@
    In addition to sometimes (critically) affecting the accuracy, mass scaling often leads to a non-diagonal scaled mass matrix, thereby partly undermining the computational savings from the reduction in the number of iterations. Surprisingly few authors have investigated this tradeoff and the time savings (if any) might come at the price of increased storage requirements. Some of the methods proposed are plainly impractical, if not completely hopeless and surely cannot compete with a lumped mass, left alone accuracy concerns. Due to the non-diagonal scaled mass, researchers have resorted to preconditioned iterative methods [45, 22] or Cholesky factorizations [46, 40] for solving linear systems. The former has also driven condition number analyses for the scaled mass.
@@
    Mass scaling methods are first and foremost designed to increase the critical time step but amazingly, step size estimates are rarely provided. Some heuristic and “analytical” step size estimates were derived for FEA [25, 26] and IGA [47, 48], often by means of characteristic lengths. However, they generally do not guarantee a stable step size [49], unless derived from rigorous bounds based e.g. on the Gershgorin circles [49, 50, 51] or on Ostrowski’s bound [52].
In view of the points raised above, our article complements previous numerical studies with a theoretical analysis, which we feel is desperately needed. Our analysis offers a unifying picture and a clearer positioning of the main methods proposed in the literature and is substantiated with some eigenvalue, step size and condition number estimates. Our study focuses exclusively on theoretical aspects and sheds light on the numerical observations reported in the literature. The few numerical experiments of this article are only meant to validate the theoretical predictions. Although mostly developed and applied for $C^{0}$ discretizations, some of the methods discussed herein are also valuable for smooth spline discretizations (i.e. IGA) and fit in the framework of outlier removal techniques provided one substitutes elements for patches.
@@ 2 Mass scaling
  • Conventional mass scaling (CMS), also sometimes called regular mass scaling, only scales the diagonal entries of the mass matrix. It is called uniform when all diagonal entries are scaled uniformly. Although it preserves the diagonal structure of the lumped mass matrix, it also scales down all eigenvalues and adversely effects the dynamics, unless performed locally on critical elements within non-uniform meshes that severely constrain the step size [24, 27, 28, 35, 25, 26]. CMS is also typically used for scaling rotational degrees of freedom of beam or shell finite element formulations [30, 31]. Despite being utterly inaccurate, this strategy is still widely used in commercial software packages, such as LS-DYNA [54]. Nevertheless, as shown in [27, 29], simply reformulating the original model may sometimes drastically improve the accuracy.
@@ 2.1.1 Preliminaries
Instead of treating $E$ as some random perturbation, we will review some of the actual strategies proposed in the literature and position them in a more general context. It turns out none of methods reviewed in this section are fundamentally new. As a matter of fact, their theoretical foundations were already laid out decades ago [56, 57, 58]. We seize this occasion to recall some fundamental, sometimes overlooked, properties of generalized eigenproblems. Throughout this article, we consider the set of pairs of matrices $V:=\mathbb{R}^{n\times n}\times\mathbb{R}^{n\times n}$, endowed with an obvious vector space structure:
@@ 2.1.2 Linear fractional transformations
Uniform mass scaling simply scales all eigenvalues by $\mu>0$ and therefore also affects the smallest ones. This disastrously impacts the dynamics, as testified by the numerical experiments in [24]. On the contrary, if the scaling parameter is wisely chosen, stiffness proportional mass scaling allows to selectively scale down the largest eigenvalues while nearly preserving the smallest ones. Olovsson et al. [1] recommend choosing $\mu=10^{k}/\lambda_{n}$ with $k\in\mathbb{N}$. If $\lambda_{n}/\lambda_{1}\gg 10^{k}$, the smallest eigenvalues are barely affected while the largest ones are roughly scaled down by a factor of $10^{k}$. In practice, computing $\mu$ only requires a (coarse) estimation for $\lambda_{n}$. Unfortunately, the pleasing property of stiffness proportional SMS comes at the price of a non-diagonal scaled mass matrix $\overline{M}=M+\mu K$. In fact, if one uses this scaled mass matrix in an explicit scheme for linear elasto-dynamics, he might as well consider using an implicit unconditionally stable Newmark method from the very start, whose workload per iteration is similar. Additionally, the method is ill-suited to nonlinear problems undergoing large rotations and deformations as the stiffness matrix changes during the course of the simulation and is typically only accessible through matrix-vector products. Stiffness proportional SMS would require reassembling the stiffness matrix at each iteration, which is prohibitively expensive and is the main reason why, according to some authors [28, 42], the method has not been implemented in commercial software packages. Nevertheless, according to Askes et al. [62], stiffness proportional SMS is employed in other branches of mechanics, albeit under different names.
@@ 2.1.3 Global deflation
Instead of transforming the entire spectrum, it is also possible to modify only a few selected eigenvalues. In numerical linear algebra, this procedure is commonly known as deflation. Its origin dates back to the early work of Hotelling [20] and is well-known for standard eigenproblems. For generalized eigenproblems, the procedure is described e.g. in [60]. For $C^{0}$ finite element discretizations, González and Park [23] revisited this idea and improved the earlier work of Tkachuk and Bischoff [22] by proposing a fast solver for the scaled mass matrix based on the Woodbury matrix identity. While the method remains wholly impractical for $C^{0}$ discretizations due to the sheer number of inaccurate frequencies, it becomes much more appealing for maximally smooth isogeometric discretizations featuring far fewer inaccurate frequencies [11, 13, 19]. It is briefly recalled here.
@@
Although the scaled mass matrix is non-diagonal, [23, Equation (28)] provides an explicit expression for its inverse, based on the Woodbury matrix identity. Unfortunately, this strategy still requires computing some of the largest eigenpairs of a (generally very large) matrix pair, which is prohibitively expensive for $C^{0}$ finite element discretizations due to the “optical branches” of the spectrum contributing too many inaccurate high frequencies. The method becomes much more affordable and even realistic for maximally smooth isogeometric discretizations featuring far fewer inaccurate high frequencies [19, Theorems A.1 and A.2] whose separation from the rest of the spectrum is also a key advantage in accelerating eigensolvers based on Krylov subspaces. Alternatively, deflation strategies may rely on low-frequency modes combined with a projector in the orthogonal complement [23]. Although such methods do not offer any computational advantages in this setting, they form the cornerstone of inexact ad hoc deflation strategies reviewed in sections 2.2.4 and 2.2.5.
@@ 2.2.1 Preliminaries
The upper bound of theorem 2.15 is usually tight [67] and translates into a relatively large and conservative step size estimate for applications in explicit dynamics. However, neither theorem 2.15 nor lemma 2.14 are applicable for bounding the eigenvalues of $CK$, where $C$ is a reciprocal mass matrix (i.e. a sparse approximate inverse of the mass matrix) [42, 43]. Indeed, $CK$ does not have the form described in eq. 2.5 and is generally not even symmetric. Instead, upper bounds based on the classical Gershgorin circle theorem [50] or its generalization given by Ostrowski’s bound [52] have been suggested. In [25], the authors have noted that the latter produced large overestimates for the element largest eigenfrequency when using $M^{-1}K$. This is not surprising given that pre-multiplication of $K$ with the diagonal matrix $M^{-1}$ will only rescale its rows. Substituting $M^{-1}K$ with the similar (and symmetric) matrix $M^{-1/2}KM^{-1/2}$ sometimes produces sharper bounds [51]. Combining the Gershgorin theorem with a diagonal similarity transformation might also improve the bounds [68, Corollary 6.1.6], particularly for beam, plate or shell problems, when the matrix entries have different physical units [52].
In contrast, several authors in the engineering community have undergone the painstaking task of deriving analytical or heuristic estimates for the largest element frequencies (see e.g. [25, 26, 47, 48]). Given the small size of element eigenvalue problems and the continuous improvement of computing capabilities, we would advise computing them numerically in parallel instead of relying on ad hoc (not necessarily conservative) estimates. It also provides a safe pathway in case element formulations are further altered due to reduced integration or locking-free formulations as well as in trimmed configurations where penalty terms are typically added to the system matrices [69, 40]. Furthermore, as noted in [47], the cost of local eigenvalue computations may be reduced by pre-selecting critical elements earmarked based on coarser but cheaper eigenvalue estimates.
The sharpness of the upper bound of theorem 2.15 also highlights how individual elements might negatively impact the critical time step and is an incentive for modifying the element formulations prior to assembly. This reasoning drove early developments of local mass scaling techniques [24]. We are now ready to analyze some of the practical schemes proposed in the literature.
@@ 3.1 Ad hoc local SMS
is the Rayleigh quotient. In this case, the perturbation scales down all eigenvalues uniformly and therefore preserves the eigenvalue ordering (i.e. $i_{k}=k$) for $k\geq 7$. These results are shown graphically in fig. 3.1 for $\beta=1$ and confirm our theoretical findings. The results are far more interesting when considering the method of Hoffmann et al. Also in this case it turns out all eigenvectors of $(K_{e},M_{e})$ are eigenvectors of $(\overline{M}_{e},M_{e})$. The engineers’ deep understanding of the deformation modes, which are eigenvectors of the stiffness matrix, certainly guided the construction of the scaling matrix. These modes bear different names and are carefully listed and grouped in [25, Figure 3]. Applying lemma 2.11 once again, we deduce that
@@
where the overline now obviously refers to scaled quantities for the method of Hoffmann et al. In this case, $Q_{e}(\mathbf{u}_{k})$ may take any value among those listed in eq. 2.13. Therefore the scaling may not preserve the eigenvalue ordering. The values of $Q_{e}(\mathbf{u}_{k})$ for $k=7,\dots,m$ are shown in fig. 3.2 for $\beta=1$ alongside the eigenvalues of the scaled and original matrix pairs. It appears that $\lambda_{i_{20}}(K_{e},\overline{M}_{e})=\lambda_{i_{21}}(K_{e},\overline{M}_{e})<\lambda_{i_{19}}(K_{e},\overline{M}_{e})$ and confirms that the scaling does not preserve the eigenvalue ordering. Generally speaking, comparing the $k$th largest eigenvalues for the original and scaled matrix pairs is only relevant for “small” perturbations. In other cases, one should better understand how the original eigenvalues are transformed.
@@
The insights drawn locally will prove useful for illustrating some global properties. As benchmark, we consider a thin metal strip of length $\times$ width $\times$ thickness dimensions $200\times 20\times 2$ mm, whose Young modulus, Poisson ratio and density are again $E=207$ GPa, $\nu=0.3$ and $\rho=7800$ kg/m3, respectively. The plate is discretized with $40\times 5\times 4$ nodes in each respective dimension, producing $348$ linear hexahedral elements. The resulting mesh is shown in fig. 3.3. In this example, we test the tightness of the bounds in corollaries 2.20 and 2.22 for the methods of Olovsson et al. [1] and Hoffmann et al. [35], respectively. The increase of the critical time step is displayed in fig. 3.4 along with its upper bound for different values of the scaling parameter $\beta$ ranging from $1$ to $500$. For the method of Olovsson et al., the upper bound is very sharp over the entire range of values tested, whereas for the method of Hoffmann et al., the upper bound is satisfactory only for moderate values of $\beta$. For larger values, the ratio flattens out and the critical time step most likely becomes constrained by other deformation modes, not targeted by the method of Hoffmann et al. Nevertheless, both methods may achieve a similar increase of the critical time step, albeit for different values of $\beta$.
@@
Following standard practice, fig. 3.5 show the ratio of scaled over unscaled frequencies for selected values of $\beta$. Note that the upper bounds of corollaries 2.20 and 2.22 (valid over the entire spectrum) become lower bounds in this case. Admittedly, except preserving perhaps the first couple of eigenfrequencies, the method of Olovsson et al. seems quite inaccurate and fig. 5(a) fuels existing concerns raised in [22, 35]. On the contrary, the method of Hoffmann et al. seems vastly superior, achieving a similar increase of the critical time step while yielding greater accuracy of the lower frequencies. Interestingly, the overall trend reported in fig. 3.5. is reminiscent of the local construction and the targeted modes.
@@
If linear systems with the scaled mass matrix are solved iteratively, as advocated in [45], its condition number is often monitored. The upper bounds in corollaries 2.20 and 2.22 suggest an increase in the number of iterations of iterative solvers (e.g. Conjugate Gradients) as the scaling parameter grows larger and may potentially offset the saving in the number of time steps. Fortunately, as shown in fig. 3.6, the upper bounds seem quite pessimistic in this case. Although the growth is indeed linear, the rate is smaller than predicted. The numerically observed rates are shown in dotted line. The reason for the discrepancy between the theoretical rates and the observed ones is quite subtle and is due to a step in the proof of lemma 2.14. Namely, the bounds $1\leq\|\mathsf{L}\mathbf{x}\|_{2}^{2}\leq p_{\max}$ may be loose for specific choices of $\mathbf{x}$. For the method of Olovsson et al., one may derive a refined bound in the asymptotic regime as $\beta\to\infty$. For this purpose, we recall that for a lumped mass matrix $M$ with uniform elements and material density, $\kappa(M)=p_{\max}$, where $p_{\max}\in\mathbb{N}^{*}$ is the maximum number of elements to which a node is connected. In our example, $p_{\max}=8$ and is attained for interior nodes. Furthermore, from the proof of lemma 2.14, the following bounds hold:
@@
$\|\mathsf{L}\mathbf{x}_{1}\|_{2}^{2}=\frac{1}{n}\sum_{e=1}^{N}\sum_{j=1}^{m}1=\frac{mN}{n},$
where $N$ is the number of elements, $m$ is the number of degrees of freedom per element and $n$ is the total number of degrees of freedom. Finally, putting all the pieces back together,

### v2

@@ 1 Introduction and background
Finite element analysis (FEA) has long established itself as an indispensable tool in many engineering disciples. Particularly in structural dynamics, one of its founding applications, where the method is extensively used for simulating the deformation and vibration of various structures and solids, including plates, shells and beams. In this context, explicit time integration is favored in several circumstances. Indeed, the dynamics of the problem often impose a natural restriction on the step size [1], and enlarging it may lead to convergence issues within implicit unconditionally stable methods, specifically for nonlinear problems requiring the solution of a nonlinear system of equations at each iteration. On the contrary, the smaller step size used within explicit methods ensures greater robustness. Moreover, explicit methods offer the possibility of “lumping” the mass matrix enabling memory-efficient matrix-free implementations. Mass lumping consists in directly substituting the mass matrix with an ad hoc diagonal approximation [2, 3, 4, 5]. Thus, explicit methods may avoid solving costly linear or nonlinear systems of equations, left alone the burden of linearizing the stiffness.
@@
$\Delta t_{c}=\frac{2}{\omega_{n}}$ (1.1)
@@
where $\omega_{n}$ is the largest frequency of the discrete system [6, 7]. The largest discrete frequencies of classical $C^{0}$ finite element methods are wholly inaccurate and fortunately rarely contribute significantly to the solution. They form the so-called “optical branches” of the spectrum and diverge with the mesh size and polynomial degree [8, 9]. Smooth isogeometric analysis (IGA) [10] features far fewer inaccurate frequencies [11, 12, 13, 14]. They usually form noticeable spikes in the upper part of the spectrum and were coined outliers for this very reason [11]. Regardless of the discretization technique and smoothness, large inaccurate eigenvalues severely restrict the critical time step and removing or dampening them is paramount in explicit dynamics. The method for removing them, however, usually depends on the discretization technique. In isogeometric analysis, so-called outlier removal techniques commonly exploit the smoothness and tensor product nature of spline spaces (see e.g. [11, 14, 15, 16, 17, 18, 19]). These methods may also preserve the consistency of the discrete formulation and therefore higher rates of convergence. In more general cases with significantly less structure, one may resort to mass scaling. Broadly speaking, mass scaling consists in dampening higher frequencies by altering the discrete formulation and, more specifically, the mass matrix. Conventional mass scaling (CMS) [20, 21], which only scales the diagonal entries of the mass matrix, is the most straightforward way of achieving this goal. Unfortunately, it also heavily deteriorates the accuracy of the lower frequencies and mode shapes, and, subsequently, the numerical solution. Selective mass scaling (SMS) is specifically designed for scaling down the highest frequencies while somewhat preserving the lowest ones. Thus, it still affects all eigenvalues, but not uniformly. Mathematically speaking, the idea is reminiscent of inexact deflation techniques, which consist in removing unwanted eigenvalues from the spectrum of a matrix. Exact deflation techniques have been known for decades and originated from the early work of Hotelling [22] for standard eigenproblems but are mostly impractical. Practical strategies, either global or local, are instead essentially heuristic and oftentimes little is known of their theoretical properties. Global strategies directly modify the assembled system matrices while local ones operate on the element matrices prior to assembly. The former are generally easier to analyze than the latter and are among the oldest methods. In the late 1990s, Macek and Aubert [23] and later Olovsson et al. [1] independently proposed the same global mass scaling procedure, based on linear fractional transformations of matrix pairs. The method was later generalized in [24] to rational polynomials, where the authors also stated optimality requirements for SMS. These methods are fully understood and transform (nonlinearly) the eigenvalues of a matrix pair while preserving its eigenvectors. Although the methods affect all eigenvalues, they may heavily damp the largest ones, while nearly preserving the smallest ones. In the same article, the authors also proposed an exact deflation method only scaling down the largest eigenvalues of the system. The idea was later revived and thoroughly improved in [25] but remains impractical, unless only a few well-separated eigenvalues must be deflated [19]. As we will see, all of these methods are merely adaptions or reformulations of well-known linear algebra facts, mostly known since decades.
@@
    In addition to sometimes (critically) affecting the accuracy, mass scaling often leads to a non-diagonal scaled mass matrix, thereby partly undermining the computational savings from the reduction in the number of iterations. Surprisingly few authors have investigated this tradeoff and the time savings (if any) might come at the price of increased storage requirements. Some of the methods proposed are impractical and may not compete with a lumped mass, left alone accuracy concerns. Due to the non-diagonal scaled mass, researchers have resorted to preconditioned iterative methods [45, 24] or Cholesky factorizations [46, 40] for solving linear systems. The former has also driven condition number analyses for the scaled mass.
@@
    Mass scaling methods are first and foremost designed to increase the critical time step but theoretical bounds on the step size are rarely provided. Some heuristic and “analytical” step size estimates were derived for FEA [27, 28] and IGA [47, 48], often by means of characteristic lengths. However, they generally do not guarantee a stable step size [49], unless derived from rigorous bounds based e.g. on the Gershgorin circles [49, 50, 51] or on Ostrowski’s bound [52].
In view of the points raised above, our article complements previous numerical studies with a theoretical analysis, which we feel is needed. Our analysis offers a unifying picture and a clearer positioning of the main methods proposed in the literature and is substantiated with some eigenvalue, step size and condition number estimates. Our study focuses exclusively on theoretical aspects and sheds light on the numerical observations reported in the literature. The few numerical experiments of this article are only meant to validate the theoretical predictions. Although mostly developed and applied for $C^{0}$ discretizations, some of the methods discussed herein are also valuable for smooth spline discretizations (i.e. IGA) and fit in the framework of outlier removal techniques provided one substitutes elements for patches.
@@ 2 Mass scaling
  • Conventional mass scaling (CMS), also sometimes called regular mass scaling, only scales the diagonal entries of the mass matrix. It is called uniform when all diagonal entries are scaled uniformly. Although it preserves the diagonal structure of the lumped mass matrix, it also scales down all eigenvalues and adversely effects the dynamics, unless performed locally on critical elements within non-uniform meshes that severely constrain the step size [26, 29, 30, 35, 27, 28]. CMS is also typically used for scaling rotational degrees of freedom of beam or shell finite element formulations [20, 21]. Despite being oftentimes inaccurate, this strategy is still widely used in commercial software packages, such as LS-DYNA [54]. Nevertheless, as shown in [29, 31], simply reformulating the original model may sometimes improve the accuracy.
@@ 2.1.1 Preliminaries
Instead of treating $E$ as some random perturbation, we will review some of the actual strategies proposed in the literature and position them in a more general context. It turns out none of methods reviewed in this section are fundamentally new. As a matter of fact, their theoretical foundations were already laid out decades ago [56, 57, 58]. We start by providing some definitions and recalling a few fundamental properties of generalized eigenproblems. Throughout this article, we consider the set of pairs of matrices $V:=\mathbb{R}^{n\times n}\times\mathbb{R}^{n\times n}$, endowed with an obvious vector space structure:
@@ 2.1.2 Linear fractional transformations
Uniform mass scaling simply scales all eigenvalues by $\mu>0$ and therefore also affects the smallest ones. This impacts the accuracy of the dynamics, as testified by the numerical experiments in [26]. On the contrary, if the scaling parameter is wisely chosen, stiffness proportional mass scaling allows to selectively scale down the largest eigenvalues while nearly preserving the smallest ones. Olovsson et al. [1] recommend choosing $\mu=10^{k}/\lambda_{n}$ with $k\in\mathbb{N}$. If $\lambda_{n}/\lambda_{1}\gg 10^{k}$, the smallest eigenvalues are barely affected while the largest ones are roughly scaled down by a factor of $10^{k}$. In practice, computing $\mu$ only requires a (coarse) estimation for $\lambda_{n}$. Unfortunately, the pleasing property of stiffness proportional SMS comes at the price of a non-diagonal scaled mass matrix $\overline{M}=M+\mu K$. In fact, if one uses this scaled mass matrix in an explicit scheme for linear elasto-dynamics, he might as well consider using an implicit unconditionally stable Newmark method from the very start, whose workload per iteration is similar. The method is even less suited to nonlinear problems undergoing large rotations and deformations as the stiffness matrix changes during the course of the simulation and is typically only accessible through matrix-vector products. Stiffness proportional SMS would require reassembling the stiffness matrix at each iteration, which is prohibitively expensive and is the main reason why, according to some authors [30, 42], the method has not been implemented in commercial software packages. Nevertheless, according to Askes et al. [62], stiffness proportional SMS is employed in other branches of mechanics, albeit under different names.
@@ 2.1.3 Global deflation
Instead of transforming the entire spectrum, it is also possible to modify only a few selected eigenvalues. In numerical linear algebra, this procedure is commonly known as deflation. Its origin dates back to the early work of Hotelling [22] and is well-known for standard eigenproblems. For generalized eigenproblems, the procedure is described e.g. in [60]. For $C^{0}$ finite element discretizations, González and Park [25] revisited this idea and improved the earlier work of Tkachuk and Bischoff [24] by proposing a fast solver for the scaled mass matrix based on the Woodbury matrix identity. While the method remains impractical for $C^{0}$ discretizations due to the sheer number of inaccurate frequencies, it becomes much more appealing for maximally smooth isogeometric discretizations featuring far fewer inaccurate frequencies [11, 13, 19]. It is briefly recalled here.
@@
Although the scaled mass matrix is non-diagonal, [25, Equation (28)] provides an explicit expression for its inverse, based on the Woodbury matrix identity. Unfortunately, this strategy still requires computing some of the largest eigenpairs of a (generally very large) matrix pair, which is prohibitively expensive for $C^{0}$ finite element discretizations due to the “optical branches” of the spectrum contributing too many inaccurate high frequencies. The method becomes much more affordable and even realistic for maximally smooth isogeometric discretizations featuring far fewer inaccurate high frequencies [19, Theorems A.1 and A.2] whose separation from the rest of the spectrum is also a key advantage in accelerating eigensolvers based on Krylov subspaces. Alternatively, deflation strategies may rely on low-frequency modes combined with a projector in the orthogonal complement [25]. Although such methods do not offer any computational advantages in this setting, they form the cornerstone of inexact ad hoc deflation strategies reviewed in Sections 2.2.4 and 2.2.5.
@@ 2.2.1 Preliminaries
The upper bound of Theorem 2.15 is usually tight [67] and translates into a relatively large and conservative step size estimate for applications in explicit dynamics. However, neither Theorem 2.15 nor Lemma 2.14 are applicable for bounding the eigenvalues of $CK$, where $C$ is a reciprocal mass matrix (i.e. a sparse approximate inverse of the mass matrix) [42, 43]. Indeed, $CK$ does not have the form described in eq. 2.5 and is generally not even symmetric. Instead, upper bounds based on the classical Gershgorin circle theorem [50] or its generalization given by Ostrowski’s bound [52] have been suggested. In [27], the authors have noted that the former produced large overestimates for the element largest eigenfrequency when using $M^{-1}K$. This is not surprising given that pre-multiplication of $K$ with the diagonal matrix $M^{-1}$ will only rescale its rows. Substituting $M^{-1}K$ with the similar (and symmetric) matrix $M^{-1/2}KM^{-1/2}$ sometimes produces sharper bounds [51]. Combining the Gershgorin theorem with a diagonal similarity transformation might also improve the bounds [68, Corollary 6.1.6], particularly for beam, plate or shell problems, when the matrix entries have different physical units [52].
In contrast, several authors in the engineering community have undertaken the task of deriving analytical or heuristic estimates for the largest element frequencies (see e.g. [27, 28, 47, 48]). However, these estimates are not necessarily conservative and rarely account for common changes to element formulations due, for instance, to reduced integration, locking-free formulations or the addition of penalty terms [69, 40]. Nevertheless, they may help identify critical elements [47], before using refined estimates.
As a matter of fact, the sharpness of the upper bound of Theorem 2.15 also highlights how individual elements might negatively impact the critical time step and is an incentive for modifying the element formulations prior to assembly. This reasoning drove early developments of local mass scaling techniques [26]. We are now ready to analyze some of the practical schemes proposed in the literature.
@@ 3.1 Ad hoc local SMS
is the Rayleigh quotient. In this case, the perturbation scales down all eigenvalues uniformly and therefore preserves the eigenvalue ordering (i.e. $i_{k}=k$) for $k\geq 7$. These results are shown graphically in Figure 3.1 for $\beta=1$ and confirm our theoretical findings. The results are far more interesting when considering the method of Hoffmann et al. Also in this case it turns out all eigenvectors of $(K_{e},M_{e})$ are eigenvectors of $(\overline{M}_{e},M_{e})$. The deep understanding of the deformation modes from engineering practice certainly guided the construction of the scaling matrix. These modes bear different names and are carefully listed and grouped in [27, Figure 3]. Applying Lemma 2.11 once again, we deduce that
@@
where the overline now obviously refers to scaled quantities for the method of Hoffmann et al. In this case, $Q_{e}(\mathbf{u}_{k})$ may take any value among those listed in eq. 2.13. Therefore, the scaling may not preserve the eigenvalue ordering. The values of $Q_{e}(\mathbf{u}_{k})$ for $k=7,\dots,m$ are shown in Figure 3.2 for $\beta=1$ alongside the eigenvalues of the scaled and original matrix pairs. It appears that $\lambda_{i_{20}}(K_{e},\overline{M}_{e})=\lambda_{i_{21}}(K_{e},\overline{M}_{e})<\lambda_{i_{19}}(K_{e},\overline{M}_{e})$ and confirms that the scaling does not preserve the eigenvalue ordering. Generally speaking, comparing the $k$th largest eigenvalues for the original and scaled matrix pairs is only relevant for “small” perturbations. In other cases, one should better understand how the original eigenvalues are transformed.
@@
The insights drawn locally will prove useful for illustrating some global properties. As benchmark, we consider a thin metal strip of length $\times$ width $\times$ thickness dimensions $200\times 20\times 2$ mm, whose Young modulus, Poisson ratio and density are again $E=207$ GPa, $\nu=0.3$ and $\rho=7800$ kg/m3, respectively. The plate is discretized with $40\times 5\times 4$ nodes in each respective dimension, producing $348$ linear hexahedral elements. The resulting mesh is shown in Figure 3.3. In this example, we test the tightness of the bounds in Corollaries 2.20 and 2.22 for the methods of Olovsson et al. [1] and Hoffmann et al. [35], respectively. The increase of the critical time step is displayed in Figure 3.4 along with its upper bound for different values of the scaling parameter $\beta$ ranging from $1$ to $500$. For the method of Olovsson et al., the upper bound is very sharp over the entire range of values tested, whereas for the method of Hoffmann et al., the upper bound is satisfactory only for moderate values of $\beta$. For larger values, the ratio flattens out and the critical time step most likely becomes constrained by other deformation modes, not targeted by the method of Hoffmann et al. Nevertheless, both methods may achieve a similar increase of the critical time step, albeit for different values of $\beta$.
@@
Following standard practice, Figure 3.5 show the ratio of scaled over unscaled frequencies for selected values of $\beta$. Note that the upper bounds of Corollaries 2.20 and 2.22 (valid over the entire spectrum) become lower bounds in this case. Admittedly, except preserving perhaps the first couple of eigenfrequencies, the method of Olovsson et al. seems quite inaccurate and Figure 5(a) fuels existing concerns raised in [24, 35]. On the contrary, the method of Hoffmann et al. seems vastly superior, achieving a similar increase of the critical time step while yielding greater accuracy of the lower frequencies. Interestingly, the overall trend reported in Figure 3.5. is reminiscent of the local construction and the targeted modes.
@@
If linear systems with the scaled mass matrix are solved iteratively, as advocated in [45], its condition number is often monitored. The upper bounds in Corollaries 2.20 and 2.22 suggest an increase in the number of iterations of iterative solvers (e.g. Conjugate Gradients) as the scaling parameter grows larger and may potentially offset the saving in the number of time steps. Fortunately, as shown in Figure 3.6, the upper bounds seem quite pessimistic in this case. Although the growth is indeed linear, the rate is smaller than predicted. The numerically observed rates are shown in dotted line. The reason for the discrepancy between the theoretical rates and the observed ones is quite subtle and is due to a step in the proof of Lemma 2.14. Namely, the bounds $1\leq\|\mathsf{L}\mathbf{x}\|_{2}^{2}\leq p_{\max}$ may be loose for specific choices of $\mathbf{x}$. For the method of Olovsson et al., one may derive a refined bound in the asymptotic regime as $\beta\to\infty$. For this purpose, we recall that for a lumped mass matrix $M$ with uniform elements and material density, $\kappa(M)=p_{\max}$, where $p_{\max}\in\mathbb{N}^{*}$ is the maximum number of elements to which a node is connected. In our example, $p_{\max}=8$ and is attained for interior nodes. Furthermore, from the proof of Lemma 2.14, the following bounds hold:
@@
$\|\mathsf{L}\mathbf{x}_{1}\|_{2}^{2}=\frac{1}{n}\sum_{i=1}^{n}p_{i}=\frac{1}{n}\sum_{e=1}^{N}\sum_{j=1}^{m}1=\frac{mN}{n},$
reduces to the average connectivity, where $N$ is the number of elements, $m$ is the number of degrees of freedom per element and $n$ is the total number of degrees of freedom. Finally, putting all the pieces back together,